# MCQG-SRefine: Multiple Choice Question Generation and Evaluation with Iterative Self-Critique, Correction, and Comparison Feedback

### v2

@@ D.1 Settings
To explore RQ3, we utilized ratings generated during the MCQG-SRefine critique step as LLM-as-Judge (rating) scores. For LLM-as-Judge (comparison), we employed the same guidelines and settings used in the human evaluation from RQ1 in main results, prompting the model to choose a preferred output between the two system-generated questions (e.g., GPT-4 and MCQG-SRefine). To ensure the results were not biased by the inherent positional preferences of the LLM (e.g., position biasWang et al. (2023a); Zheng et al. (2023); Zeng et al. (2023), we collected preferences from two sequence settings: one where the GPT-4 generated output was shown first followed by the MCQG-SRefine output, and the reverse order. We filtered out preferences that did not match across the two sequence settings, using the remaining data for further analysis. To ensure the results were not influenced by length bias (i.e., LLMs like GPT-4 prefer longer generations during their automatic evaluation)  Wang et al. (2023c); Zeng et al. (2023), we deliberately included those data points in the evaluation data which had the least context length ratio for the initial and the MCQG-SRefine questions. These examples constituted half the human evaluation data and the rest were chosen randomly from the dataset. Notably, the guidelines used in human evaluation, LLM-as-Judge (rating), and LLM-as-Judge (comparison) were consistent. Thus, by standardizing the evaluation data, we could fairly compare the judgments of Expert X and LLM-as-Judge (rating/comparison) under identical settings to assess the reliability of LLM-as-Judge.
@@ D.2 Results
For RQ3, we found that the rating method performs slightly better than the comparison method. Percentage Agreement between GPT-4-Comparison and Expert X: Total agreement: 53.8% (Human: 58.33%, Machine: 50%). Percentage Agreement between GPT-4-Rating and Expert X: Total agreement: 61.538% (Human: 58.33%, Machine: 64.285%). For RQ4, as shown in Appendix Table 9 and  18, we observed that GPT-4 as the LLM-as-judge has the least position bias (35% of the data showed inconsistent results under two different orders, while other models showed more than 50% inconsistency). Moreover, GPT-4 has a relatively higher correlation with human evaluations in the valid data (i.e., data without position bias) with a percentage agreement of 61.53% and a Cohen’s kappa of 0.2121. On the other hand, Claude-3-opus showed 52.5% position bias but had the highest correlation with human evaluation in the valid data, with a percentage agreement of 68.42% and a Cohen’s kappa of 0.3152.

### v4

@@ D.1 Settings
To explore RQ3, we utilized ratings generated during the MCQG-SRefine critique step as LLM-as-Judge (rating) scores. For LLM-as-Judge (comparison), we employed the same guidelines and settings used in the human evaluation from RQ1 in main results, prompting the model to choose a preferred output between the two system-generated questions (e.g., GPT-4 and MCQG-SRefine). To ensure the results were not biased by the inherent positional preferences of the LLM (e.g., position biasWang et al. (2023b); Zheng et al. (2023); Zeng et al. (2023), we collected preferences from two sequence settings: one where the GPT-4 generated output was shown first followed by the MCQG-SRefine output, and the reverse order. We filtered out preferences that did not match across the two sequence settings, using the remaining data for further analysis. To ensure the results were not influenced by length bias (i.e., LLMs like GPT-4 prefer longer generations during their automatic evaluation)  Wang et al. (2023d); Zeng et al. (2023), we deliberately included those data points in the evaluation data which had the least context length ratio for the initial and the MCQG-SRefine questions. These examples constituted half the human evaluation data and the rest were chosen randomly from the dataset. Notably, the guidelines used in human evaluation, LLM-as-Judge (rating), and LLM-as-Judge (comparison) were consistent. Thus, by standardizing the evaluation data, we could fairly compare the judgments of Expert X and LLM-as-Judge (rating/comparison) under identical settings to assess the reliability of LLM-as-Judge.
@@ D.2 Results
For RQ3, we found that the rating method performs slightly better than the comparison method. Percentage Agreement between GPT-4-Comparison and Expert X: Total agreement: 53.8% (Human: 58.33%, Machine: 50%). Percentage Agreement between GPT-4-Rating and Expert X: Total agreement: 61.538% (Human: 58.33%, Machine: 64.285%). For RQ4, as shown in Appendix Table 9 and  19, we observed that GPT-4 as the LLM-as-judge has the least position bias (35% of the data showed inconsistent results under two different orders, while other models showed more than 50% inconsistency). Moreover, GPT-4 has a relatively higher correlation with human evaluations in the valid data (i.e., data without position bias) with a percentage agreement of 61.53% and a Cohen’s kappa of 0.2121. On the other hand, Claude-3-opus showed 52.5% position bias but had the highest correlation with human evaluation in the valid data, with a percentage agreement of 68.42% and a Cohen’s kappa of 0.3152.
@@
Finally, we also calculated self-BLEU scores as diversity metrics in two scenarios: 1. Self-BLEU for GPT-4 generated USMLE MCQs. 2. Self-BLEU for USMLE MCQs generated by the MCQG-SRefine framework. Our analysis shows that MCQG-SRefine reduces the self-BLEU score by approximately 25.9% compared to GPT-4’s original generation (lower self-BLEU indicates less similarity and hence greater diversity). We believe this is because SRefine adjusts the initial contexts—typically more homogeneous due to covering broader portions of the clinical note—to make them more focused on individual topics or test points, thereby enhancing diversity. Below are the results:
Model Human Machine
GPT-4 0.5839 0.5793
MCQG-SRefine 0.4326 0.4322
Table 12: Performance comparison of GPT-4 and GPT-4 + MCQG-SRefine on Human Data and Machine Data.